OpenStudy (anonymous):

Use the price-demand equation below to find the values of p for which demand is elastic and the values for which demand is inelastic. Assume that price and demand are both positive. x=f(p)=330-30p

OpenStudy (helder_edwin):

the elasticity is given by \[\large \varepsilon=\frac{\partial\ln x}{\partial\ln p}=\frac{-30p}{300-30p} \]

OpenStudy (helder_edwin):

to see where the demand is elastic solve \[\large \varepsilon>1 \] to see where the demand is inelastic solve \[\large \varepsilon<1 \]

OpenStudy (anonymous):

How would I set up the equation to see where the demand would be either elastic or inelastic?